# Data-driven decision-making under uncertainty with entropic risk measure

### v1

@@ Appendix B Additional results
$\ \leq\ \frac{1}{u^{*}}\left(\hat{u}_{N}({{\boldsymbol{z}}})-u({{\boldsymbol{z}}})\right),$
where the second inequality follows from $u(z)\geq u^{*}$. The above inequality implies that
$\left\{\mathop{\rm sup}_{{{\boldsymbol{z}}}\in\mathcal{Z}}\left|\hat{\rho}_{N}({{\boldsymbol{z}}})-\rho({{\boldsymbol{z}}})\right|\geq\gamma\right\}\subseteq\left\{\mathop{\rm sup}_{{{\boldsymbol{z}}}\in\mathcal{Z}}\left|\hat{u}_{N}({{\boldsymbol{z}}})-u({{\boldsymbol{z}}})\right|\geq\alpha\gamma u^{*}\right\}$
$\Rightarrow{\mathbb{P}}\left(\mathop{\rm sup}_{{{\boldsymbol{z}}}\in\mathcal{Z}}\left|\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))-\rho^{*})\right|\geq\gamma\right)\leq{\mathbb{P}}\left(\mathop{\rm sup}_{{{\boldsymbol{z}}}\in\mathcal{Z}}\left|\hat{u}_{N}({{\boldsymbol{z}}})-u({{\boldsymbol{z}}})\right|\geq\alpha\gamma u^{*}\right).$
On combining with (53), we obtain
@@
${\mathbb{P}}\left\{\mathop{\rm sup}_{{{\boldsymbol{z}}}\in\mathcal{Z}}\left|\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))-\rho^{*})\right|\geq\gamma\right\}\leq\frac{A}{N^{\frac{p}{2}}(\alpha\gamma u^{*})^{p}}.$ (54)
Finally, we note that
$\lvert\rho_{\text{SAA}}-\rho^{*}\rvert$ $\ =\ \left|\mathop{\rm min}_{{{\boldsymbol{z}}}\in\mathcal{Z}}\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))-\mathop{\rm min}_{z\in\mathcal{Z}}\rho_{{\mathbb{P}}}(\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))\right|$
$\ \leq\ \mathop{\rm max}_{{{\boldsymbol{z}}}\in\mathcal{Z}}\left|\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))-\rho_{{\mathbb{P}}}(\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))\right|.$
On combining the above inequality with (54) and substituting $u^{*}=\exp(\alpha\rho^{*})$, we obtain
${\mathbb{P}}\left(\lvert\rho_{\text{SAA}}-\rho^{*}\rvert\geq\gamma\right)\leq\frac{A}{N^{\frac{p}{2}}(\gamma\alpha\exp(\alpha\rho^{*}))^{p}}.$
Theorem 5
Suppose $\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})$ is Lipschitz in ${{\boldsymbol{\xi}}}$ for all ${{\boldsymbol{z}}}$ with Lipschitz constant $L$, and assume the SAA solution’s entropic risk converges to the true risk at rate $\text{O}(N^{-\frac{p}{2}})$ for sufficiently large $N$. If the radius of the ambiguity set is $cN^{-\frac{p}{2}}$, then the risk $\rho_{\text{DRO}}$ converges to the optimal risk $\rho^{*}$ in probability, i.e., for any $\gamma>0$, there exists a constant $A>0$ such that:
${\mathbb{P}}(\lvert\rho_{\mbox{\scriptsize DRO}}-\rho^{*}\rvert\geq\gamma)\leq\frac{A}{N^{\frac{p}{2}}(\alpha\left(\gamma-LcN^{-\frac{p}{2}}\right)\exp(\alpha\rho^{*}))^{p}}$
Proof. Since ^$\hat{{\mathbb{P}}}_{N}\in\mathcal{B}_{\infty}(\epsilon)$, the following inequality holds for all ${{\boldsymbol{z}}}$
$\mathop{\rm sup}_{\mathbb{Q}\in\mathcal{B}_{\infty}(\epsilon)}\frac{1}{\alpha}\log[{\mathbb{E}}_{\mathbb{Q}}(\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})))]\geq\frac{1}{\alpha}\log({\mathbb{E}}_{\hat{{\mathbb{P}}}_{N}}[\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})))].$
Taking the minimum over ${{\boldsymbol{z}}}\in\mathcal{Z}$ on both sides of the above inequality, we obtain
$\mathop{\rm min}_{z\in\mathcal{Z}}\mathop{\rm sup}_{\mathbb{Q}\in\mathcal{B}_{\infty}(\epsilon)}\frac{1}{\alpha}\log[{\mathbb{E}}_{\mathbb{Q}}(\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})))]\geq\mathop{\rm min}_{z\in\mathcal{Z}}\frac{1}{\alpha}\log({\mathbb{E}}_{\hat{{\mathbb{P}}}_{N}}[\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})))]$
@@
To prove the other side of the inequality, we will use the lipschitz continuity of $\ell$ in ${{\boldsymbol{\xi}}}$ for all ${{\boldsymbol{z}}}\in\mathcal{Z}$:
$|\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}_{1})-\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}_{2})|\leq L\|{{\boldsymbol{\xi}}}_{1}-{{\boldsymbol{\xi}}}_{2}\|.$
where $L$ is the lipschitz constant. From Bertsimas et al., (2023), the following problems are equivalent
$\mathop{\rm sup}_{\mathbb{Q}\in\mathcal{B}_{\infty}(\epsilon)}\frac{1}{\alpha}\log[{\mathbb{E}}_{\mathbb{Q}}(\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})))]=\frac{1}{\alpha}\log[{\mathbb{E}}_{\hat{{\mathbb{P}}}_{N}}\mathop{\rm sup}_{\hat{{{\boldsymbol{\xi}}}}_{i}:\|{{\boldsymbol{\xi}}}-\hat{{{\boldsymbol{\xi}}}}_{i}\|\leq\epsilon}\exp(\alpha\ell({{\boldsymbol{z}}},\hat{{{\boldsymbol{\xi}}}}_{i}))]$ (55)
Then, it follows from the lipschitz continuity of $\ell$ and $\|\hat{{{\boldsymbol{\xi}}}}_{i}-\hat{{{\boldsymbol{\xi}}}}_{i}\|\leq\epsilon$ that
$|\ell({{\boldsymbol{z}}},\hat{{{\boldsymbol{\xi}}}}_{i})-\ell({{\boldsymbol{z}}},\hat{{{\boldsymbol{\xi}}}}_{i})|\leq L\|\hat{{{\boldsymbol{\xi}}}}_{i}-\hat{{{\boldsymbol{\xi}}}}_{i}\|\leq L\epsilon\Rightarrow\ell({{\boldsymbol{z}}},\hat{{{\boldsymbol{\xi}}}}_{i})\leq\ell({{\boldsymbol{z}}},\hat{{{\boldsymbol{\xi}}}}_{i})+L\epsilon$
The above inequality implies that
$\mathop{\rm sup}_{\hat{{{\boldsymbol{\xi}}}}_{i}:\|{{\boldsymbol{\xi}}}-\hat{{{\boldsymbol{\xi}}}}_{i}\|\leq\epsilon}\exp(\alpha\ell({{\boldsymbol{z}}},\hat{{{\boldsymbol{\xi}}}}_{i}))\leq\exp(\alpha(\ell({{\boldsymbol{z}}},\hat{{{\boldsymbol{\xi}}}}_{i})+L\epsilon))$
Substituting in (55), we obtain:
$\mathop{\rm sup}_{\mathbb{Q}\in\mathcal{B}_{\infty}(\epsilon)}\frac{1}{\alpha}\log[{\mathbb{E}}_{\mathbb{Q}}(\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})))]\leq\frac{1}{\alpha}\log[{\mathbb{E}}_{\hat{{\mathbb{P}}}_{N}}(\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})+\alpha L\epsilon))]$
$\Rightarrow\mathop{\rm sup}_{\mathbb{Q}\in\mathcal{B}_{\infty}(\epsilon)}\frac{1}{\alpha}\log[{\mathbb{E}}_{\mathbb{Q}}(\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})))]\leq\frac{1}{\alpha}\log[{\mathbb{E}}_{\hat{{\mathbb{P}}}_{N}}(\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))])+L\epsilon$
From the above inequality, we can show that
$\rho_{\mbox{\scriptsize DRO}}=\rho_{\mbox{\scriptsize SAA}}+\delta L\epsilon,$
where $0\leq\delta\leq 1$. Taking $\epsilon=cN^{-\frac{p}{2}}$ with $c>0$, we obtain,
$\rho_{\mbox{\scriptsize DRO}}=\rho_{\mbox{\scriptsize SAA}}+\delta LcN^{-\frac{p}{2}}.$
Using the triangle inequality, we have:
$\lvert\rho_{\mbox{\scriptsize DRO}}-\rho^{*}\rvert\leq\lvert\rho_{\mbox{\scriptsize DRO}}-\rho_{\mbox{\scriptsize SAA}}\rvert+\lvert\rho_{\mbox{\scriptsize SAA}}-\rho^{*}\rvert.$
Therefore for $\gamma>0$, we have
$\left\{\lvert\rho_{\mbox{\scriptsize DRO}}-\rho^{*}\rvert\geq\gamma\right\}\ \subseteq\ \left\{\rho_{\mbox{\scriptsize DRO}}-\rho_{\mbox{\scriptsize SAA}}\rvert+\lvert\rho_{\mbox{\scriptsize SAA}}-\rho^{*}\rvert\geq\gamma\right\}$
$\Rightarrow{\mathbb{P}}(\lvert\rho_{\mbox{\scriptsize DRO}}-\rho^{*}\rvert\geq\gamma)\leq{\mathbb{P}}(\lvert\rho_{\mbox{\scriptsize DRO}}-\rho_{\mbox{\scriptsize SAA}}\rvert+\lvert\rho_{\mbox{\scriptsize SAA}}-\rho^{*}\rvert\geq\gamma)={\mathbb{P}}(\delta LcN^{-\frac{p}{2}}+\lvert\rho_{\mbox{\scriptsize SAA}}-\rho^{*}\rvert\geq\gamma^{\prime}).$
Combining the above inequality with Lemma 2, we obtain
${\mathbb{P}}(\lvert\rho_{\mbox{\scriptsize DRO}}-\rho^{*}\rvert\geq\gamma)\leq{\mathbb{P}}\left\{\lvert\rho_{\mbox{\scriptsize SAA}}-\rho^{*}\rvert\geq\gamma-\delta LcN^{-\frac{p}{2}}\right\}\leq\frac{A}{N^{\frac{p}{2}}(\alpha\left(\gamma-\delta LcN^{-\frac{p}{2}}\right)\exp(\alpha\rho^{*}))^{p}}.$
Since $\delta\in[0,1]$, we obtain
${\mathbb{P}}(\lvert\rho_{\mbox{\scriptsize DRO}}-\rho^{*}\rvert\geq\gamma)\leq\frac{A}{N^{\frac{p}{2}}\left(\alpha\left(\gamma-LcN^{-\frac{p}{2}}\right)\exp(\alpha\rho^{*})\right)^{p}}.$
Theorem 6
$p$-Wasserstein DRO with entropic risk measure results in unbounded loss if $p<\infty$.
Proof. Let $u({{\boldsymbol{z}}})$ denote the worst-case expected utility for a given decision ${{\boldsymbol{z}}}$:
$u({{\boldsymbol{z}}}):=\mathop{\rm sup}_{\mathbb{Q}\in\mathcal{B}_{p}(\epsilon)}u_{\mathbb{Q}}(\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})):=\mathbb{E}_{\mathbb{Q}}[\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))].$ (56)
From Gao and Kleywegt, (Lemma 2, Proposition 2, 2023), the worst-case disutility $u({{\boldsymbol{z}}})$ is infinite with $p$-Wasserstein $(p<\infty)$ ambiguity set since the loss function $g({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})=\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))$ does not satisfy the growth condition: $g({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})\leq L\|{{\boldsymbol{\xi}}}-{{\boldsymbol{\xi}}}_{0}\|^{p}+M$, for all ${{\boldsymbol{\xi}}}\in\Xi$ and some constants $L>0,M>0$, and ${{\boldsymbol{\xi}}}_{0}\in\Xi$ for any ${{\boldsymbol{z}}}\in\mathcal{Z}$. Next, we will prove that $u({{\boldsymbol{z}}})$ is unbounded for all ${{\boldsymbol{z}}}\in\mathcal{Z}$ implies that $\rho({{\boldsymbol{z}}})$ is unbounded for all ${{\boldsymbol{z}}}\in\mathcal{Z}$ by contradiction. Suppose that $\rho({{\boldsymbol{z}}})$ is bounded. Then $\exists\,M<\infty$ such that
$\mathop{\rm sup}_{\mathbb{Q}\in\mathcal{B}_{p}(\epsilon)}\frac{1}{\alpha}\log\left(u({{\boldsymbol{z}}})\right)\leq M.$ (57)
Since the above relation holds for all $\mathbb{Q}$ and $\log$ is a monotonic function, we obtain:
$u_{\mathbb{Q}}(\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))\leq\exp(\alpha M)\text{ for all }\mathbb{Q}\in\mathcal{B}_{p}.$
This implies that $u({{\boldsymbol{z}}})$ is bounded, which contradicts the fact that $u({{\boldsymbol{z}}})$ is unbounded.
B.1 Influence function
A way to explain the underestimation effect of a finite sample can be seen when looking at the loss distribution through the influence function. The influence function measures the sensitivity of entropic risk to small changes in the data. The influence function of a statistic $T$ at a point $\xi$ for a distribution ${\mathbb{P}}$ is given by
$\text{IF}(\xi;T,{\mathbb{P}})=\lim_{\varepsilon\to 0}\frac{T((1-\varepsilon){\mathbb{P}}+\varepsilon\delta_{\xi})-T({\mathbb{P}})}{\varepsilon},$
where $\delta_{\xi}$ is the Dirac distribution at the point $\xi$. The influence function of entropic risk measure is given by
$\text{IF}(\xi;\rho,{\mathbb{P}})=$ $\lim_{\varepsilon\to 0^{+}}\frac{\log\left((1-\varepsilon){\mathbb{E}}_{\mathbb{P}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))]+\varepsilon\exp(\alpha\ell({{\boldsymbol{\xi}}}))\right)-\log\left({\mathbb{E}}_{\mathbb{P}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))]\right)}{\alpha\varepsilon}$
$=\lim_{\varepsilon\to 0^{+}}\frac{\log\left((1-\varepsilon)+\varepsilon\frac{\exp(\alpha\ell({{\boldsymbol{\xi}}}))}{{\mathbb{E}}_{\mathbb{P}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))]}\right)}{\alpha\varepsilon}=-\frac{1}{\alpha}+\frac{\exp(\alpha\ell({{\boldsymbol{\xi}}}))}{\alpha{\mathbb{E}}_{\mathbb{P}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))]}.$
For entropic risk measure $\rho$, $\text{IF}(\hat{\xi},\rho,{\mathbb{P}})$ for any support point $\hat{\xi}$ can be obtained in closed form for $\Gamma$-distributed loss $\xi$ as follows:
$\text{IF}(\hat{\xi};\rho,{\mathbb{P}})=-\frac{1}{\alpha}+\frac{\exp(\alpha\hat{\xi})}{\alpha\left(1-\lambda\alpha\right)^{-\kappa}}.$
In Figure 10, we plot the histogram of $1000$ loss values generated from the $\Gamma(10,0.45)$ distribution. We compute the average IF over samples in each bin and darker shades represent higher (normalized) values of average IF. It can be seen that tail events have highest impact on the entropic risk but occur with very low probability. Therefore, these high impact scenarios are likely not to be included in a finite sample, resulting in the underestimation effect.
Figure 10: Histogram of the $1000$ loss samples from a $\Gamma(10,0.45)$ distribution and the average IF is computed by taking the mean of IF over scenarios in each bin with the darker blue shade representing higher (normalized) average IF.
B.2 Bias correction using OIC
From Theorem 1 in Iyengar et al., (2023), it follows that
${\mathbb{E}}_{\mathcal{D}_{N}}{\mathbb{E}}_{{\mathbb{P}}}[h(\hat{\theta},\ell({{\boldsymbol{\xi}}}))]={\mathbb{E}}_{\mathcal{D}_{N}}[h(\hat{\theta},\ell(\hat{{{\boldsymbol{\xi}}}}_{i}))]-\frac{1}{N}{\mathbb{E}}_{{\mathbb{P}}}[\nabla_{\theta}h(\hat{\theta},\ell({{\boldsymbol{\xi}}}))\text{IF}(\ell({{\boldsymbol{\xi}}}),\rho,{\mathbb{P}})]+o\left(\frac{1}{N}\right),$
where the influence function is given by $\text{IF}(\ell({{\boldsymbol{\xi}}}),\rho,{\mathbb{P}})=-\frac{1}{\alpha}+\frac{\exp(\alpha\ell(\xi))}{\alpha{\mathbb{E}}_{{\mathbb{P}}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))]}.$ The bias of a decision $\hat{\theta}$ is given by $\delta_{\mbox{\scriptsize OIC}}=-\frac{1}{N}{\mathbb{E}}_{{\mathbb{P}}}[\nabla_{\theta}h(\hat{\theta},\ell({{\boldsymbol{\xi}}}))\text{IF}(\ell({{\boldsymbol{\xi}}}),\rho,{\mathbb{P}})].$ When ${\mathbb{P}}$ is not known, Iyengar et al., (2023) propose an approximate bias correction term by replacing the true distribution with ^$\hat{{\mathbb{P}}}_{N}$. On substituting $\nabla_{\theta}h(\theta,\ell({{\boldsymbol{\xi}}}))=1-\exp(\alpha(\ell({{\boldsymbol{\xi}}})-\theta))$:
$\hat{\delta}_{\text{OIC}}$ $\ =\ \frac{1}{N^{2}}\sum_{i=1}^{N}\left(1-\exp(\alpha(\ell(\hat{{{\boldsymbol{\xi}}}}_{i})-\hat{\theta}))\right)\left(\frac{1}{\alpha}-\frac{\exp(\alpha\ell(\hat{{{\boldsymbol{\xi}}}}_{i}))}{\alpha\sum_{j\in[N]}\frac{\exp(\alpha\ell(\hat{{{\boldsymbol{\xi}}}}_{j}))}{N}}\right)$
$\ =\ \frac{1}{\alpha N^{2}}\sum_{i=1}^{N}\left(1-\exp(\alpha(\ell(\hat{{{\boldsymbol{\xi}}}}_{i})-\hat{\theta}))\right)^{2}.$
B.3 Differential sampling from GMM

### v3

@@ B.2 Bias mitigation using OIC
${\mathbb{E}}[{\mathbb{E}}_{{\mathbb{P}}}(h(\hat{t},\ell({{\boldsymbol{\xi}}})))]={\mathbb{E}}[h(\hat{t},\ell(\hat{{{\boldsymbol{\xi}}}}_{i}))]\underbrace{-\frac{1}{N}{\mathbb{E}}_{{\mathbb{P}}}[\nabla_{t}h(t^{*},\ell({{\boldsymbol{\xi}}}))\text{IF}(t^{*})]}_{\delta_{\texttt{OIC}}}+o\left(\frac{1}{N}\right),$
where expectation is with respect to the randomness of $\mathcal{D}_{N}$. For $h(t,\ell({{\boldsymbol{\xi}}}))=t+\frac{1}{\alpha}\exp(\alpha(\ell({{\boldsymbol{\xi}}})-t))-\frac{1}{\alpha}$, we know that $t^{*}=\rho_{{\mathbb{P}}}(\ell({{\boldsymbol{\xi}}}))={\mathbb{E}}_{{\mathbb{P}}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))]$, $\nabla_{t}h(t^{*},\ell({{\boldsymbol{\xi}}}))=1-\exp(\alpha(\ell({{\boldsymbol{\xi}}})-t^{*}))$ and $\nabla_{t,t}h(t^{*},\ell({{\boldsymbol{\xi}}}))=\alpha\exp(\alpha(\ell({{\boldsymbol{\xi}}})-t^{*}))$. The influence function in the expression of $\delta_{\texttt{OIC}}$ is obtained as follows:
@@
$\text{IF}(t^{*})\,=\,-\left({\mathbb{E}}_{\mathbb{P}}[\nabla^{2}_{t,t}h(t^{*},\ell({{\boldsymbol{\xi}}}))]\right)^{-1}\nabla_{t}h(t^{*},\ell({{\boldsymbol{\xi}}}))\,=\,-\frac{1-\exp(\alpha(\ell({{\boldsymbol{\xi}}})-t^{*}))}{{\mathbb{E}}_{{\mathbb{P}}}[\alpha\exp(\alpha(\ell({{\boldsymbol{\xi}}})-t^{*}))]}$
Next, we substitute the value of $\text{IF}(t^{*})$ and $\nabla_{t}h(t^{*},\ell({{\boldsymbol{\xi}}}))$ to obtain the bias of a decision $t^{*}$:
$\delta_{\texttt{OIC}}\,=$ $-\frac{1}{N}{\mathbb{E}}_{{\mathbb{P}}}[\nabla_{t}h(t^{*},\ell({{\boldsymbol{\xi}}}))\text{IF}(t^{*})]$
@@
[…]
Since ${\mathbb{P}}$ is not known, Iyengar et al., (2023) replace ${\mathbb{P}}$ with $\hat{{\mathbb{P}}}$ to obtain their estimator, $\rho_{\texttt{OIC}}:=\hat{t}+\text{Var}_{\hat{{\mathbb{P}}}_{N}}(\exp(\alpha\ell({{\boldsymbol{\xi}}})))/(N\alpha(\mathbb{E}_{{\hat{{\mathbb{P}}}_{N}}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))])^{2})$.
B.3 Fitting a GMM
Entropic risk matching.
After each update of the parameters of a GMM using the gradient descent procedure described in Algorithm 2, the parameters are projected back into the feasible region for a valid GMM. This ensures that the mixing weights $\hat{\boldsymbol{\pi}}_{t+1}$ remain valid probabilities (which is achieved using a softmax function), and that the standard deviations $\hat{\boldsymbol{\sigma}}_{t+1}$ are positive (enforced by taking the maximum of $\exp(-5)$ and $\hat{\boldsymbol{\sigma}}_{t+1}$). The number of components $Y$ in the GMM is selected by CV based on the Wasserstein distance between the distribution of the entropic risk of samples drawn from fitted GMM and the distribution of entropic risk constructed from the scenarios in the validation set. In all numerical experiments, we set the maximum iterations $T=30000$ and tolerance $\epsilon=\exp(-9)$.
Matching the extremes.
The distribution of the maxima of $n$ i.i.d samples from a normal distribution $\mathcal{N}(\mu,\sigma)$ is given by $\left(\Phi_{\mu,\sigma}\right)^{n}$ where $\Phi_{\mu,\sigma}$ is the cdf of a normally distributed random variable with mean $\mu$ and $\sigma$. We find the parameters of the normal distribution by matching the 50th and 90th quantiles of $F_{\mathcal{M}}(.)$ to the corresponding quantiles of $\left(\Phi_{\mu,\sigma}\right)^{n}$:
$\mu+\sigma\Phi_{0,1}^{-1}(0.5^{1/n})=F_{\mathcal{M}}(0.5)$
$\mu+\sigma\Phi_{0,1}^{-1}(0.9^{1/n})=F_{\mathcal{M}}(0.9),$
B.4 Differential sampling from GMM
In Algorithm 2, differentiable samples are generated using Algorithm 6. This algorithm leverages the reparameterization trick (Kingma et al.,, 2015) for continuous distributions and the Gumbel-Softmax trick (Jang et al.,, 2017; Maddison et al.,, 2017) for discrete distributions. The Gumbel-Softmax trick allows approximate, differentiable sampling of mixture components, while Gaussian samples are obtained by combining deterministic transformations of the parameters with random noise. As a result, gradients can flow through both the discrete and continuous sampling steps. This enables the minimization of the Wasserstein distance between the empirical and model-based entropic risk distributions.